# Interdisciplinary Digital Twin Engine InterTwin for calorimeter simulation

### v1

@@ 2 Choosing a model for fast calorimeter simulation
The interTwin project is a European Union–funded initiative that aims to develop a cross-disciplinary Digital Twin Engine (DTE), providing an open-source modular platform for the creation and deployment of scientific digital twins. Its architecture combines core infrastructure modules with domain-specific thematic modules. A central component of this ecosystem is itwinai, a Python-based framework designed to streamline the use of artificial intelligence in digital-twin applications. It offers experiment tracking and model registry through MLflow, support for distributed training across HPC and cloud environments, and workflow definition via configuration files. These features reduce the engineering burden associated with machine-learning workflows.
Among the many scientific use cases, the high-energy physics application focuses on accelerating calorimeter simulations through generative models. Within this use case, two models are available. The first is 3DGAN [6], a generative adversarial network that produces three-dimensional calorimeter showers represented as voxelized energy depositions in calorimeters. Its generator is conditioned on the incident particle energy and the angle of incidence. By exploiting three-dimensional convolutional architectures, 3DGAN achieves speed-ups of several orders of magnitude over Geant4 while preserving key calorimetric observables such as total energy deposition and longitudinal and lateral shower profiles. The second model, CaloINN, has also been integrated into the interTwin framework. Its implementation within itwinai benefits from the framework’s experiment tracking, configuration management, and workflow orchestration, thereby ensuring reproducibility and facilitating large-scale experiments. The presence of both 3DGAN and CaloINN in the interTwin ecosystem demonstrates the feasibility of applying the platform to high-energy physics simulations.
@@ 4 Fixing distribution tails
The central objective of the ongoing research is to improve the fidelity of the most promising generative model, CaloINN. In practice, this requires achieving accuracies better than 10% in the distribution tails at the $10^{-3}$ quantile level. Achieving this level of precision is essential for ensuring that fast simulation tools do not bias rare-event searches or precision measurements. Without any corrections, the performance of the CaloINN model is shown on Figure 1.
@@
Building on this idea, our ongoing work explores fine-tuning procedures at the inference stage. CaloINN is first to be trained on datasets with artificially enhanced tails, thereby over-representing extreme calorimeter showers. A dedicated classifier is then to be introduced to distinguish between generative model outputs and Geant4-based simulations, enabling the estimation of density ratios across the feature space. These ratios are subsequently applied during inference to reweight and correct the generated distributions, effectively pushing the tails toward the Geant4 reference. This approach draws inspiration from variance-reduction techniques [8] and is designed to preserve the overall speed advantage of generative models. The additional computational cost is expected to remain below a factor of two, which is acceptable given that the baseline speed-up relative to Geant4 typically spans two to three orders of magnitude.

### v2

@@ 2 Choosing a model for fast calorimeter simulation
The interTwin project is a European Union–funded initiative that aims to develop a cross-disciplinary Digital Twin Engine (DTE), providing an open-source modular platform for the creation and deployment of scientific digital twins. Its architecture combines core infrastructure modules with domain-specific thematic modules. A central component of this ecosystem is itwinai, a Python-based framework designed to streamline the use of artificial intelligence in digital-twin applications. It offers experiment tracking and model registry through MLFlow [6], support for distributed training across HPC and cloud environments, and workflow definition via configuration files. These features reduce the engineering burden associated with machine-learning workflows.
Among the many scientific use cases, the high-energy physics application focuses on accelerating calorimeter simulations through generative models. Within this use case, two models are available. The first is 3DGAN [7], a generative adversarial network that produces three-dimensional calorimeter showers represented as voxelized energy depositions in calorimeters. Its generator is conditioned on the incident particle energy and the angle of incidence. By exploiting three-dimensional convolutional architectures, 3DGAN achieves speed-ups of several orders of magnitude over Geant4 while preserving key calorimetric observables such as total energy deposition and longitudinal and lateral shower profiles. The second model, CaloINN, has also been integrated into the interTwin framework. Within itwinai, the implementation benefits from configuration management that consolidates all data, model, and training parameters into a single configuration file. This includes access and setup of experiment tracking via MLFlow and distributed computing strategies. Such approach ensures reproducibility and facilitates large-scale experiments. The presence of both 3DGAN and CaloINN in the interTwin ecosystem demonstrates the feasibility of applying the platform to high-energy physics simulations.
@@ 4 Fixing distribution tails
The central objective of the ongoing research is to improve the fidelity of the CaloINN model. In practice, this requires achieving accuracies better than 10% in the distribution tails at the $10^{-3}$ quantile level. Achieving this level of precision is essential for ensuring that fast simulation tools do not bias rare-event searches or precision measurements. Without any corrections, the performance of the CaloINN model is shown on Figure 1.
@@
Building on this idea, our ongoing work explores fine-tuning procedures at the inference stage. CaloINN is first to be trained on datasets with artificially enhanced tails, thereby over-representing extreme calorimeter showers. Such showers can be selected based on anomaly scores computed using Isolation Forest [9] or Local Outlier Factor [10] algorithms, which quantify how strongly a given shower deviates from the rest of the dataset. A dedicated classifier is then to be introduced to distinguish between generative model outputs and Geant4-based simulations, enabling the estimation of density ratios across the feature space. These ratios are subsequently applied during inference to reweight and correct the generated distributions, pushing the tails toward the Geant4 reference. The concrete implementation of this procedure is currently under investigation. This approach draws inspiration from variance-reduction techniques [11] and is designed to preserve the overall speed advantage of generative models. The additional computational cost is expected to remain below a factor of two, which is acceptable given that the baseline speed-up relative to Geant4 typically spans two to three orders of magnitude.